If Marichu can finish typing a report in 1 hour and 30 minutes, what part of the report can she finish in 1 hour?
Answer by stanbon(75887)

If Marichu can finish typing a report in 1 hour and 30 minutes, what part of the report can she finish in 1 hour?
---------------
Note: 1 hr and 30 minutes = 90 minutes
-----------------
1/1.5 = x/1
x = 1/(3/2)
x = 2/3 of a report in 1 hour.
============
Cheers,
Stan H.